OpenStudy (anonymous):

describe how to transform the graph of f into the graph of g. f(x) = x4 and g(x) = -(- x)4

OpenStudy (jdoe0001):

\(\bf f(x) = x^4 \qquad g(x) = -(- x)^4\\\quad \\ (-x)^4 \implies (-x)\times (-x)\times (-x)\times (-x) \implies x\\ \quad \\ g(x) = -(- x)^4 \implies g(x) = -x^4 \)

OpenStudy (jdoe0001):

\(\bf from\qquad f(x) = x^4 \qquad to \qquad g(x) = - x^4\) so what do you think?

OpenStudy (debbieg):

@jdoe0001 I think you meant to say in the 2nd line: \(\Large (-x)^4 \implies (-x)\times (-x)\times (-x)\times (-x) \implies x\color{red}{^4}\)

OpenStudy (jdoe0001):

ohh yes, thanks =)

OpenStudy (jdoe0001):

\(\bf \bf f(x) = x^4 \qquad g(x) = -(- x)^4\\\quad \\ (-x)^4 \implies (-x)\times (-x)\times (-x)\times (-x) \implies x^4\\ \quad \\ g(x) = -(- x)^4 \implies g(x) = -x^4\)

OpenStudy (jdoe0001):

hehe

OpenStudy (anonymous):

Thanks. What about this one: describe how to transform the graph of f into the graph of g. f(x) = x4 and g(x) = -(- x4)

OpenStudy (anonymous):

Thanks. What about this one: describe how to transform the graph of f into the graph of g. f(x) = x4 and g(x) = -(- x4)